OpenStudy (anonymous):

Cyclopropane, an anesthetic, decomposes to propane.The reaction is first order with respect to cyclopropane with k = 9.2 s-1 at 1000°C. a) Calculate the half-life of the reaction. b) How much of a 3.00 g sample of cyclopropane remains after 0.50 sec at 1000°C?

OpenStudy (jfraser):

Minor correction, but cyclopropane can't decompose into propAne, but it can decompose into propEne:  The formulas must be the same, so C3H6(cyclo) --> C3H6 (linear) a) plug into the equation for determining half life when given k:\[t_{\frac{1}{2}} = \frac{\ln(2)}{k}\] b) plug the time, k, and mass of cyclopropane into the rate law equation for a first-order reaction and solve for A_t: \[A_t = A_0*e^{-kt}\]